# Global well-posedness of the elastic-viscous-plastic sea-ice model with the inviscid Voigt-regularisation

### v1

@@ 1.3 Main result of this paper
$\partial_{t}u=\nabla\cdot\sigma+\mathcal{T}_{a}+\mathcal{T}_{w}+\Omega u^{\perp}-g\nabla H_{0},$ (1.12a)
$\dfrac{1}{\mathcal{E}}\partial_{t}\bigl{(}\sigma-\alpha^{2}\Delta\sigma\bigr{)}+\dfrac{4{\mathcal{D}}}{P}(\sigma-\frac{1}{2}\operatorname{Tr}\sigma\mathbb{I}_{2})+\dfrac{{\mathcal{D}}}{2P}\operatorname{Tr}\sigma\mathbb{I}_{2}+\dfrac{{\mathcal{D}}}{2}\mathbb{I}_{2}=D(u),$ (1.12b)
$u\lvert_{t=0}=u_{0},\quad\sigma\lvert_{t=0}=\sigma_{0}.$ (1.12c)
@@
Let $u_{0}\in H^{2}(\mathbb{T}^{2})$ and $\sigma_{0}\in H^{3}(\mathbb{T}^{2})$ (which is symmetric) be given. Moreover, we assume that $P>0$ is a given constant, $U_{a},U_{w},H_{0}\in C([0,T];H^{4}(\mathbb{T}^{2}))$ and we let $\theta\in\left[0,\frac{\pi}{4}\right]$. The Voigt-EVP system (1.12) has a unique global strong solution $(u,\sigma)$ on the time interval $[0,T]$ for any $T>0$. The solution $(u,\sigma)$ has the following regularity
$u\in C([0,T];H^{2}(\mathbb{T}^{2})),\quad\sigma\in C([0,T];H^{3}(\mathbb{T}^{2})).$
In addition, we have the estimate
$\begin{gathered}\lVert u\rVert_{C([0,T];H^{2}(\mathbb{T}^{2}))}+\lVert\sigma\rVert_{C([0,T];H^{3}(\mathbb{T}^{2}))}+\lVert\partial_{t}u\rVert_{L^{\infty}((0,T);H^{2}(\mathbb{T}^{2}))}+\lVert\partial_{t}\sigma\rVert_{L^{\infty}((0,T);H^{3}(\mathbb{T}^{2}))}\\ \leq C\exp(\exp(\exp(\exp(\exp(CT))))),\end{gathered}$
@@
We note that the superexponential bound from Theorem 1.2 originates from the repeated use of (logarithmic) Grönwall inequalities for the a priori estimates at the $L^{2}(\mathbb{T}^{2})$, $\dot{H}^{1}(\mathbb{T}^{2})$ and $\dot{H}^{2}(\mathbb{T}^{2})$ level.
@@
    The Voigt regularisation in equation (1.12b) is in the same spirit as the original numerical regularisation of the Hibler model from [35]. Namely, we seek an inviscid regularisation of the EVP model, which is suitable for numerical computations. The advantage of the Voigt regularisation is that on the one hand, like the original EVP model, it shares the same steady state(s) with the Hibler model. On the other hand, the Voigt regularisation constitutes a relatively modest (nonviscous) modification of the (artificial) elastic component of the rheology in the EVP model.
@@
    For the momentum equation of the Voigt-EVP model, we have omitted the advection term, in which we follow the approach in [35]. As it has been remarked in [35], the advection term is at least one order of magnitude smaller than the acceleration term and therefore it has not been included in the EVP model. In the evolution equation for the stress tensor we also do not include an advection term, as it would affect the steady states of the EVP model, which would then lose its purpose as an approximation of the Hibler model.
@@
    We note that the assumption $\theta\in\left[0,\frac{\pi}{4}\right]$ on the water turning angle, lies within the standard range in sea-ice modelling (see for example [31, 49]).
@@
$\partial_{t}u=\nabla\cdot\sigma+\mathcal{T}_{a}+\mathcal{T}_{w}+\Omega u^{\perp}-g\nabla H_{0},$ (1.13a)
$\dfrac{1}{\mathcal{E}}\partial_{t}\bigl{(}\sigma-\alpha^{2}\Delta\sigma\bigr{)}+\dfrac{4{\mathcal{D}_{\epsilon}}}{P}(\sigma-\frac{1}{2}\operatorname{Tr}\sigma\mathbb{I}_{2})+\dfrac{{\mathcal{D}_{\epsilon}}}{2P}\operatorname{Tr}\sigma\mathbb{I}_{2}+\dfrac{{\mathcal{D}_{\epsilon}}}{2}\mathbb{I}_{2}=D(u),$ (1.13b)
$u\lvert_{t=0}=u_{0},\quad\sigma\lvert_{t=0}=\sigma_{0}.$ (1.13c)
In this model, the strain rate $\mathcal{D}$ in system (1.12) has been regularised by
$\mathcal{D}_{\epsilon}\coloneqq\sqrt{\lvert D(u)\rvert^{2}+\epsilon^{2}},$ (1.14)
for some $\epsilon>0$, where we recall that $D(u)$ denotes the symmetric part of the gradient, see equation (1.1).
Then, by establishing uniform-in-$\epsilon$ estimates of the solutions to system (1.13) globally in time, we will be able to send $\epsilon\rightarrow 0$ and obtain the (unique) solution to system (1.12). As a byproduct of this proof, we also obtain the global well-posedness of the intermediate system (1.13a)-(1.13c) as stated below:
@@
Under the same conditions on $u_{0},\sigma_{0},U_{a},U_{w},H_{0},$ and $\theta$ as in Theorem 1.2, for any $T>0$ the intermediate system (1.13) is globally well-posed. The unique global solution $(u,\sigma)$ has the same regularity as in Theorem 1.2.
@@
    The benefit of introducing $\mathcal{D}_{\epsilon}$ in the intermediate system (1.13) is that it regularises the strain rate $\mathcal{D}$ and allows one to take its derivatives directly. We emphasise that such a regularisation serves as a cutoff of the bulk and shear viscosities and has an important role in sea-ice modelling as it describes the yield curve of sea-ice, as part of the viscous-plastic rheology introduced in [32].
    Moreover, it is a natural numerical regularisation for the strain rate, which was first introduced in [41], cf. [57]. Note that in practice, $\epsilon$ should be smaller than $10^{-9}\;$s-1 [7]. For these reasons the study of the intermediate system (1.13) has its value independent of its role as an approximate system for the Voigt-EVP model.
@@
    The proof of Theorem 1.2 implies that the limit as $\epsilon\rightarrow 0$ of the intermediate system (1.13), i.e. removing the bulk and shear viscosity cutoff, is stable in the sense that the solutions remain close to each other for all small $\epsilon$. Our results are therefore very much in agreement with the numerical simulations in [7], where it was found that the dynamics of the EVP model is not very much affected by the value of $\epsilon$.
@@
    In [53], the approximation (1.14) is also used for the Hibler model (1.3b) as a regularisation of the strain rate, and it is shown that for any fixed $\epsilon>0$, the resulting system of equations is locally well-posed. However, as demonstrated by the numerical evidence in [7], the limit as $\epsilon\rightarrow 0$ for the Hibler model is very unstable. The analysis of the limit as $\epsilon\rightarrow 0$ remains open.
@@
    Instead of (1.14), our analysis works for a larger class of regularisations for the strain rate $\mathcal{D}$. See Section 5 for more discussion.
@@ 1.4 Overview of the literature
Sea ice is a highly complex material and modelling its dynamics and rheology poses significant challenges. In fact, numerous sea-ice models have been developed over time [25]. The Arctic Ice Dynamics Joint Experiment suggested an elastic-plastic rheology for modelling sea ice [22]. The Hibler sea-ice model was introduced in [32], which uses a viscous-plastic rheology which has become one of the standard sea-ice rheologies.
@@
The Hibler model has been a very successful model in reproducing observed sea-ice drift [54]. However, its use in practical computations turns out to be very expensive [7, 40], particularly so when using an explicit numerical method [38]. Instead, implicit methods are employed in order to compute solutions of the Hibler model, for example the line relaxation method [68] as well as the Jacobian-free Newton Krylov solver [46, 55, 62] to mention some.
@@
The elastic-viscous-plastic (EVP) sea-ice model was introduced in [35] in order to improve the computational efficiency of sea-ice modelling. Specifically, the EVP model is a modification of the Hibler model, as an elastic (relaxation) term is introduced into the constitutive relation. The advantage of the EVP formulation over the Hibler model is that it is expected to be more amenable to the use of explicit numerical schemes for computational modelling, and hence facilitates the use of parallel computing [34], see also [36]. As a result, numerically solving the EVP model significantly reduces the computational cost compared to the use of implicit schemes for the Hibler model [7].
@@
In addition, the EVP model (1.3a) and (1.3c) has the advantage that its computational performance is not affected by the value of the strain rate regularisation parameter $\epsilon$, while the Hibler model requires a small value of $\epsilon$ in the viscosity regularisation to accurately model plastic behaviour as shown in [7], cf. Remark 1.5. Because of the aforementioned reasons the EVP model has become one of the standard and most commonly used sea-ice models in contemporary climate modelling [7, 25, 40].
@@
Over time, several modifications of the EVP model were introduced which have better numerical performance in simulations. For example, in [34] the elastic parameter was redefined to depend on the strain rate. The reason for this modification was to prevent the computed stresses from lying outside the yield curve (cf. [32]) during the (numerical) subcycling process.
@@
However, although the EVP model was originally introduced as a reformulation of the Hibler model which was more suitable for numerical purposes, it remains unclear whether it successfully approximates the Hibler model, i.e. by sending $\mathcal{E}\rightarrow\infty$ in equation (1.3c) to obtain (1.3b). In particular, it was found in [46, 54, 56, 39], see also [6], that the numerically computed solutions of the EVP and the Hibler models can differ substantially, even for very small subcycling time steps. Note that this does not necessarily imply that either sea-ice model is physically more accurate, but just that their rheologies are fundamentally different from a computational point of view [46].
@@
Nevertheless, within the subcycling time steps, the effect of the artificial elasticity in the rheology is ‘damped’ over the larger timescale, although it was found in [46] that the damping effect is weaker in areas of rigid ice, i.e. the regions with high concentration of ice and small strain rates.
@@
The first results in the rigorous analysis of sea-ice dynamics only appeared very recently. The local well-posedness of the Hibler model was proved independently in [53, 10]. Both of these papers slightly modified the original viscosity cutoff from [32] by introducing a regularisation of the form (1.14), which is required to avoid the singularity when $D(u)=0$.
@@
To be more precise, in [10], additional artificial diffusion terms were considered in the thermodynamics, which make the system fully parabolic. These additional diffusion terms made it possible to prove global existence of strong solutions for small data in [10]. Meanwhile, the study in [53] retained the original mixed hyperbolic-parabolic character of the Hibler model. A different regularisation of the viscosities was studied in [17], namely the replacement of the Heaviside cutoff by a hyperbolic tangent, where the authors establish the linear well-posedness.
@@
The interaction between sea-ice and a rigid body for the fully parabolic Hibler model was then studied in [5]. Time periodic solutions for the parabolic regularisation of the Hibler model, were constructed in [11] in the case of periodic wind forcing. See [4] for some results on the coupling with oceanic and atmospheric dynamics. The free-boundary problem of shallow (land) ice-sheets was studied in [60].
@@
In addition, in [57], under the assumption of the existence of smooth solutions, a formal $H^{1}$-estimate was derived for the Hibler model as well as for a modified version of the EVP model in the formulation introduced in [34]. This formulation of the EVP model from [34] has a very different structure to our setting. The gradient estimates in [57] strongly rely on the cutoff of the bulk and shear viscosities in the constitutive relation, and also on the linearisation of the ocean drag term. In the revised EVP case, the additional assumption that $\operatorname{Tr}\sigma=0$ is made in [57]. The a priori estimates in [57] do not seem sufficient to deduce the existence of solutions for the revised EVP model, but instead are used in [57] to check the consistency of a numerical discretisation scheme for these sea-ice models.
@@
We also note that the EVP model (1.3a) and (1.3c) shares some structural similarities with several models for non-Newtonian flows, for example the Oldroyd-B model, see for example [29, 51, 18, 42, 50, 19, 24, 21] and references therein, and other viscoelastic flow models, which have been studied in [15, 2, 14] and references therein. Finally, we note that our treatment of the viscosity regularisation (1.14) is in some sense related to results for non-Newtonian fluid models with shear dependent viscosities, see for example [3] and references therein.
@@
As has been mentioned in point $(i)$ in Remark 1.3, in this work we have chosen to use the (inviscid) Voigt regularisation for the EVP model. To the best of our knowledge, the Voigt regularisation in the context of the Navier-Stokes equations has first been rigorously studied in [58, 59]. Much later, the existence of solutions for such Kelvin-Voigt-type systems with non-constant density has also been investigated in [1] and see references therein. Moreover, the Voigt regularisation has been investigated extensively in the context of turbulence modelling, for example in [44, 43, 45]. More recently, it has also been used as a regularisation mechanism to study steady states [20], see also [37]. We also remark that employing the Voigt regularisation typically leads to equations of pseudo-parabolic type, which have been studied in [65, 63, 64] (and see references therein).
@@
The rest of this paper is organised as follows. In Section 2, we will recall some inequalities that we will use throughout this paper. In Section 3, we will establish the global well-posedness of the intermediate system (1.13). Following this, in Section 4, we finish the proof of Theorem 1.2, namely the global well-posedness of the Voigt-EVP model (1.12). We will provide some concluding remarks in Section 5.

### v2

@@ 1.3 The main result of this paper
$\partial_{t}u=\nabla\cdot\sigma+\mathcal{T}_{a}+\mathcal{T}_{w}+\Omega u^{\perp}-g\nabla H_{0},$ (12a)
$\dfrac{1}{\mathcal{E}}\partial_{t}\bigl(\sigma-\alpha^{2}\Delta\sigma\bigr)+\dfrac{4{\mathcal{D}}}{P}(\sigma-\frac{1}{2}\operatorname{Tr}\sigma\mathbb{I}_{2})+\dfrac{{\mathcal{D}}}{2P}\operatorname{Tr}\sigma\mathbb{I}_{2}+\dfrac{{\mathcal{D}}}{2}\mathbb{I}_{2}=D(u),$ (12b)
$u\lvert_{t=0}=u_{0},\quad\sigma\lvert_{t=0}=\sigma_{0}.$ (12c)
@@
Let $u_{0}\in H^{2}(\mathbb{T}^{2})$ and $\sigma_{0}\in H^{3}(\mathbb{T}^{2})$ (which is symmetric) be given. Moreover, we assume that $P>0$ is a given constant, $U_{a},U_{w},H_{0}\in C([0,T];H^{4}(\mathbb{T}^{2}))$ and we let $\theta\in\left[0,\frac{\pi}{4}\right]$. The Voigt-EVP system (12) has a unique global strong solution $(u,\sigma)$ on the time interval $[0,T]$ for any $T>0$. The solution $(u,\sigma)$ enjoys the following regularity
$u\in C([0,T];H^{2}(\mathbb{T}^{2})),\quad\sigma\in C([0,T];H^{3}(\mathbb{T}^{2})),$
and $\sigma$ is symmetric. In addition, we have the estimate
$\begin{gathered}\lVert u\rVert_{C([0,T];H^{2}(\mathbb{T}^{2}))}+\lVert\sigma\rVert_{C([0,T];H^{3}(\mathbb{T}^{2}))}+\lVert\partial_{t}u\rVert_{L^{\infty}((0,T);H^{2}(\mathbb{T}^{2}))}+\lVert\partial_{t}\sigma\rVert_{L^{\infty}((0,T);H^{3}(\mathbb{T}^{2}))}\\ \leq C\exp(\exp(\exp(CT))),\end{gathered}$
@@
We note that the superexponential bound from Theorem 1.2 originates from the repeated use of (logarithmic) Grönwall inequalities for the a priori estimates for the $L^{2}(\mathbb{T}^{2})$, $\dot{H}^{1}(\mathbb{T}^{2})$ and $\dot{H}^{2}(\mathbb{T}^{2})$ norms.
@@
    The Voigt regularisation in Eq. (12b) is in the same spirit as the original numerical regularisation of the Hibler model from Ref. [39]. Namely, we seek an inviscid regularisation of the EVP model, which is suitable for numerical computations. The advantage of the Voigt regularisation is that on the one hand, like the original EVP model, it shares the same steady state(s) with the Hibler model. On the other hand, the Voigt regularisation constitutes a relatively modest (nonviscous) modification of the (artificial) elastic component of the rheology in the EVP model.
@@
    For the momentum equation of the Voigt-EVP model, we have omitted the advection term, in which we follow the approach in Ref. [39]. As it has been remarked in Ref. [39], the advection term is at least one order of magnitude smaller than the acceleration term and therefore it has not been included in the EVP model. In the evolution equation for the stress tensor we also do not include an advection term, as it would affect the steady states of the EVP model, which would then lose its purpose as an approximation of the Hibler model.
@@
    We note that the assumption $\theta\in\left[0,\frac{\pi}{4}\right]$ on the water turning angle, lies within the standard range in sea-ice modelling (see for example Refs. [35, 53]).
@@
$\partial_{t}u=\nabla\cdot\sigma+\mathcal{T}_{a}+\mathcal{T}_{w}+\Omega u^{\perp}-g\nabla H_{0},$ (13a)
$\dfrac{1}{\mathcal{E}}\partial_{t}\bigl(\sigma-\alpha^{2}\Delta\sigma\bigr)+\dfrac{4{\mathcal{D}_{\epsilon}}}{P}(\sigma-\frac{1}{2}\operatorname{Tr}\sigma\mathbb{I}_{2})+\dfrac{{\mathcal{D}_{\epsilon}}}{2P}\operatorname{Tr}\sigma\mathbb{I}_{2}+\dfrac{{\mathcal{D}_{\epsilon}}}{2}\mathbb{I}_{2}=D(u),$ (13b)
$u\lvert_{t=0}=u_{0},\quad\sigma\lvert_{t=0}=\sigma_{0}.$ (13c)
In this model, the strain rate $\mathcal{D}$ in system (12) has been regularised by
$\mathcal{D}_{\epsilon}\coloneqq\sqrt{\lvert D(u)\rvert^{2}+\epsilon^{2}},$ (14)
for some $\epsilon>0$, where we recall that $D(u)$ denotes the symmetric part of the gradient, see Eq. (1).
Then, by establishing uniform-in-$\epsilon$ estimates of the solutions to system (13) globally in time, we will be able to send $\epsilon\rightarrow 0$ and obtain the (unique) solution to system (12). As a byproduct of this proof, we also obtain the global well-posedness of the intermediate system (13a)-(13c) as stated below:
@@
Under the same conditions on $u_{0},\sigma_{0},U_{a},U_{w},H_{0},$ and $\theta$ as in Theorem 1.2, for any $T>0$ the intermediate system (13) is globally well-posed. The unique global solution $(u,\sigma)$ has the same regularity as in Theorem 1.2.
@@
    The benefit of introducing $\mathcal{D}_{\epsilon}$ in the intermediate system (13) is that it regularises the strain rate $\mathcal{D}$ and allows one to take its derivatives directly. We emphasise that such a regularisation serves as a cutoff of the bulk and shear viscosities and has an important role in sea-ice modelling as it describes the yield curve of sea ice, as part of the viscous-plastic rheology introduced in Ref. [36].
    Moreover, it is a natural numerical regularisation for the strain rate, which was first introduced in Ref. [45], cf. Ref. [61]. Note that in practice, $\epsilon$ should be smaller than $10^{-9}\;$s-1.[7] For these reasons the study of the intermediate system (13) has its value independent of its role as an approximate system for the Voigt-EVP model.
@@
    The proof of Theorem 1.2 implies that the limit as $\epsilon\rightarrow 0$ of the intermediate system (13), i.e. removing the bulk and shear viscosity cutoff, is stable in the sense that the solutions remain close to each other for all small $\epsilon$. Our results are therefore very much in agreement with the numerical simulations in Ref. [7], where it was found that the dynamics of the EVP model is not very much affected by the value of $\epsilon$.
@@
    In Ref. [57], the approximation (14) is also used for the Hibler model (3b) as a regularisation of the strain rate, and it is shown that for any fixed $\epsilon>0$, the resulting system of equations is locally well-posed. However, as demonstrated by the numerical evidence in Ref. [7], the limit as $\epsilon\rightarrow 0$ for the Hibler model is very unstable. The analysis of the limit as $\epsilon\rightarrow 0$ remains open.
@@
    Instead of (14), our analysis works for a larger class of regularisations for the strain rate $\mathcal{D}$. See Section 5 for more discussion.
@@ 1.4 Overview of the literature
Sea ice is a highly complex material and modelling its dynamics and rheology poses significant challenges. In fact, numerous sea-ice models have been developed over time.[29] The Arctic Ice Dynamics Joint Experiment suggested an elastic-plastic rheology for modelling sea ice.[24] The Hibler sea-ice model was introduced in Ref. [36], which uses a viscous-plastic rheology which has become one of the standard sea-ice rheologies.
@@
The Hibler model has been a very successful model in reproducing observed sea-ice drift.[58] However, its use in practical computations turns out to be very expensive,[7, 44] particularly so when using an explicit numerical method.[42] Instead, implicit methods are employed in order to compute solutions of the Hibler model, for example the line relaxation method (see Ref. [72]) as well as the Jacobian-free Newton Krylov solver (see Refs. [50, 59, 66]) to mention some.
@@
The elastic-viscous-plastic (EVP) sea-ice model was introduced in Ref. [39] in order to improve the computational efficiency of sea-ice modelling. Specifically, the EVP model is a modification of the Hibler model, as an elastic (relaxation) term is introduced into the constitutive relation. The advantage of the EVP formulation over the Hibler model is that it is expected to be more amenable to the use of explicit numerical schemes for computational modelling, and hence facilitates the use of parallel computing,[38] see also Ref. [40]. As a result, numerically solving the EVP model significantly reduces the computational cost compared to the use of implicit schemes for the Hibler model.[7]
@@
In addition, the EVP model (3a) and (3c) has the advantage that its computational performance is not affected by the value of the strain rate regularisation parameter $\epsilon$, while the Hibler model requires a small value of $\epsilon$ in the viscosity regularisation to accurately model plastic behaviour as shown in Ref. [7], cf. Remark 1.5. Because of the aforementioned reasons the EVP model has become one of the standard and most commonly used sea-ice models in contemporary climate modelling.[7, 29, 44]
@@
Over time, several modifications of the EVP model were introduced which have better numerical performance in simulations. For example, in Ref. [38] the elastic parameter was redefined to depend on the strain rate. The reason for this modification was to prevent the computed stresses from lying outside the yield curve (cf. Ref. [36]) during the (numerical) subcycling process.
@@
However, although the EVP model was originally introduced as a reformulation of the Hibler model which was more suitable for numerical purposes, it remains unclear whether it successfully approximates the Hibler model, i.e. by sending $\mathcal{E}\rightarrow\infty$ in Eq. (3c) to obtain (3b). In particular, it was found in Refs. [43, 50, 58, 60], see also Ref. [6], that the numerically computed solutions of the EVP and the Hibler models can differ substantially, even for very small subcycling time steps. Note that this does not necessarily imply that either sea-ice model is physically more accurate, but just that their rheologies are fundamentally different from a computational point of view.[50]
@@
Nevertheless, within the subcycling time steps, the effect of the artificial elasticity in the rheology is ‘damped’ over the larger timescale, although it was found in Ref. [50] that the damping effect is weaker in areas of rigid ice, i.e. the regions with high concentration of ice and small strain rates. Finally, we note that several other sea-ice models have been developed in the literature. For example, in order to describe sea-ice dynamics across a range of different length scales, a multiscale model was recently developed in Ref. [26] (see also Ref. [20]).
@@
The first results in the rigorous analysis of sea-ice dynamics only appeared very recently. The local well-posedness of the Hibler model was proved independently in Refs. [57, 11]. Both of these papers slightly modified the original viscosity cutoff from Ref. [36] by introducing a regularisation of the form (14), which is required to avoid the singularity when $D(u)=0$.
@@
To be more precise, in Ref. [11], additional artificial diffusion terms were considered in the thermodynamics, which make the system fully parabolic. These additional diffusion terms made it possible to prove global existence of strong solutions for small data in Ref. [11]. Meanwhile, the study in Ref. [57] retained the original mixed hyperbolic-parabolic character of the Hibler model. Later on, the same result as in Ref. [57] was also proved in Ref. [10] with different techniques. A different regularisation of the viscosities was studied in Ref. [18], namely the replacement of the Heaviside cutoff by a hyperbolic tangent, where the authors establish the linear well-posedness (see also Ref. [17] for a study of the two-dimensional case).
@@
The interaction between sea ice and a rigid body for the fully parabolic regularisation of the Hibler model was then studied in Ref. [5]. See Ref. [4] for some results on the coupling with oceanic and atmospheric dynamics. The free-boundary problem of shallow (land) ice-sheets was studied in Ref. [64].
@@
In addition, in Ref. [61], under the assumption of the existence of smooth solutions, a formal $H^{1}$-estimate was derived for the Hibler model as well as for a modified version of the EVP model in the formulation introduced in Ref. [38]. This formulation of the EVP model from Ref. [38] has a very different structure to our setting. The gradient estimates in Ref. [61] strongly rely on the cutoff of the bulk and shear viscosities in the constitutive relation, and also on the linearisation of the ocean drag term. In the revised EVP case, the additional assumption that $\operatorname{Tr}\sigma=0$ is made in Ref. [61]. The a priori estimates in Ref. [61] do not seem sufficient to deduce the existence of solutions for the revised EVP model, but instead are used in Ref. [61] to check the consistency of a numerical discretisation scheme for these sea-ice models.
@@
We also note that the EVP model (3a) and (3c) shares some structural similarities with several models for non-Newtonian flows, for example the Oldroyd-B model, see for example Refs. [19, 21, 23, 27, 33, 46, 54, 55] and references therein, and other viscoelastic flow models, which have been studied in Refs. [2, 14, 15] and references therein. Finally, we note that our treatment of the viscosity regularisation (14) is in some sense related to results for non-Newtonian fluid models with shear dependent viscosities, see for example Ref. [3] and references therein.
@@
As has been mentioned in point $(i)$ in Remark 1.3, in this work we have chosen to use the (inviscid) Voigt regularisation for the EVP model. To the best of our knowledge, the Voigt regularisation in the context of the Navier-Stokes equations has first been rigorously studied in Refs. [62, 63]. Much later, the existence of solutions for such Kelvin-Voigt-type systems with non-constant density has also been investigated in Ref. [1] and see references therein. Moreover, the Voigt regularisation has been investigated extensively in the context of turbulence modelling, for example in Refs. [47, 48, 49]. More recently, it has also been used as a regularisation mechanism to study steady states,[22] see also Ref. [41]. We also remark that employing the Voigt regularisation typically leads to equations of pseudo-parabolic type, which have been studied in Refs. [67, 68, 69] (and see references therein).
@@
The rest of this paper is organised as follows. In Section 2, we will recall some inequalities that we will use throughout this paper. In Section 3, we will establish the global well-posedness of the intermediate system (13). Following this, in Section 4, we finish the proof of Theorem 1.2, namely the global well-posedness of the Voigt-EVP model (12). We will provide some concluding remarks in Section 5.